# Contrastive Learning on Multimodal Analysis of Electronic Health Records

### v1

@@ 4 Simulation Studies
In this section, we assess CLAIME’s performance and compare it to CL and Concate via simulation studies. When patient-level data is not available, these three methods can be deployed by performing SVD on different aggregated PMI matrices introduced in Section 2. Since patient-level data is also available in the synthesized datasets, we can additionally apply the gradient descent-based multimodal contrastive learning method (CLAIME-GD) with target function at (2.3), and the gradient descent-based contrastive learning method (CL-GD) with target function at (2.6), respectively. The suffix “-GD” is added to distinguish them from their SVD-based counterparts. To enhance training efficiency, only $10$ negative samples per patient are stochastically sampled. The penalty parameter $\lambda$ is set to $1$, while the learning rate starts at $10^{-4}$ and decays by a factor of $10$ every $10$ epochs to ensure convergence. Both gradient-based algorithms converge when their loss functions change by less than $10^{-6}$.
We generate patient embedding vector $\mathbf{c}_{i}\in\mathbb{R}^{p}$ for each patient $i$ from a multivariate normal distribution $N(\mathbf{0},\mathbf{I}_{p})$ independently. We construct the code and CUI embedding matrices $\mathbf{V}_{1}^{\star}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}\in\mathbb{R}^{d_{2}\times p}$, respectively, by generating each entry from a standard normal distribution. The row means of $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ are standardized to zero, and their maximum singular values are standardized to $1$. Then, we follow the data generation process in (2.1) to generate $T_{i}^{(1)}$ number of codes and $T_{i}^{(2)}$ number of CUIs for the $i$-th patient, where $T_{i}^{(1)},T_{i}^{(2)}$ follow a Poisson distribution with the mean of $50$ and only non-zero values are retained. Two different cases of error covariance structures are considered.
@@
For Case 1, we set $\mathbf{\Sigma}^{(M)}=\text{diag}(\mathbf{I}_{p/2},\mathbf{0})/c\in\mathbb{R}^{d_{M}\times d_{M}}$ for $M=1$ or $2$, where $c>0$ is a constant controlling the signal-to-noise ratio (SNR) of the data generating process. For Case 2, we generated $\sigma_{w}\overset{i.i.d.}{\sim}\text{Unif}(0,1)$ and set ${\bm{\Sigma}}_{w,w^{\prime}}^{(M)}=\rho^{|w-w^{\prime}|}\sigma_{w}\sigma_{w^{\prime}}/2$ for $w,w^{\prime}\in[d_{M}]$ and $M\in\{1,2\}$, where $\rho\in(0,1)$ controls both the structure and the SNR of the data generating process. Specifically, a larger $\rho$ leads to a greater signal-to-noise ratio in the synthesized data.
@@
For both cases, three experiments are separately conducted to compare the performance of CLAIME, CLAIME-GD, CL, CL-GD, and Concate under different sample sizes $n$, feature sizes $d$, and SNR. Their performance is evaluated using the metric $\text{Err}(\widehat{\mathbf{V}},\mathbf{V}^{\star})=\max_{M=1,2}\|\widehat{\mathbf{V}}_{M}\widehat{\mathbf{V}}_{M}^{\top}-\mathbf{V}_{M}^{\star}\mathbf{V}_{M}^{\star\top}\|_{\mathrm{F}}$, which measures the distance between the estimated and true embedding subspaces. We first fix the SNR by setting $c=0.2$ for Case 1 and $\rho=0.8$ for Case 2. In the first experiment, we vary the sample sizes $n$ in $\{2\times 10^{4},4\times 10^{4},6\times 10^{4},8\times 10^{4},10^{5}\}$, while fixing $d=2d_{1}=2d_{2}=1000$ and $p=4$. In the second experiment, we vary $d$ in $\{200,400,600,800,1000\}$ and set $d_{1}=d_{2}=d/2$, while fixing $n=10^{5}$ and $p=4$. For the third experiment, we first fix $n=10^{5}$, $d=2d_{1}=2d_{2}=1000$ and $p=4$. Then, we vary $c$ equally spaced between $0.2$ and $1$ for Case 1, and we vary $\rho$ equally spaced between 00 and $0.9$ for Case 2.
@@
The error metrics $\text{Err}(\widehat{\mathbf{V}},\mathbf{V}^{\star})$, averaged over $500$ repetitions, are presented in Figure 1. The results for CL and CL-GD, CLAIME and CLAIME-GD are very close to each other, as indicated by the proximity of their lines in Figures 1 (a)–(c). This suggests that employing gradient-based methods on patient-level data (i.e., CL-GD and CLAIME-GD) is comparable to utilizing SVD on summary-level PMI matrices (i.e., CL and CLAIME). This observation aligns with the theoretical findings in Propositions 2.1 and 2.2. On one hand, in Figures 1(a) and (b), the errors for CLAIME and CLAIME-GD decrease with an increase in $n$ or a decrease in $d$. This behavior is anticipated and can be theoretically inferred from the upper bounds provided in Theorems 3.13.3. In contrast, the errors for Concate, CL, and CL-GD remain unchanged. This can be theoretically justified by their lower bounds being irrelevant to both $n$ and $d$, as shown in Theorem 3.6. From the top panel of Figure 1(c), it can be observed that all methods exhibit larger errors when $c$ decreases, or in other words, the SNR decreases. Comparing to the other methods, CLAIME and CLAIME-GD demonstrate superior performance in the small SNR regions where $c\leq 0.5$. Similarly, the bottom panel of Figure 1(c) shows a consistent increase in errors across all methods with higher values of $\rho$. CLAIME and CLAIME-GD again significantly outperform the other methods when $\rho$ is larger than $0.5$. This supports our initial motivation that the proposed method will be more robust in cases when the data is noisier.
@@ 5 Application to Electronic Health Records Studies
We consider training a joint representation of codified and narrative CUI features using summary EHR data of patients with at least 1 diagnostic code of rhematoid arthritis (RA) at Mass General Brigham (MGB). The MGB RA EHR cohort consist of 53,716 patients whose longitudinal EHR data have been summarized as co-occurrence counts of EHR concepts within a $30$-day window. We include all EHR diagnosis, medication, and procedure codes which have been rolled up to higher concept levels: diagnostic codes to PheCodes [^1] https://phewascatalog.org/phecodes, procedure codes to clinical classification system (CCS) [^2] https://www.hcup-us.ahrq.gov/toolssoftware/ccs_svcsproc/ccssvcproc.jsp, medication codes to ingredient level RxNorm codes (Liu et al.,, 2005). We include all EHR codified features along with CUIs that have occurred more than $15$ times in the cohort. This results in $d=4,668$ features, including $d_{1}=3,477$ codified features with $1,776$ PheCodes for diseases, $238$ CCS codes for procedures, and $1,463$ RxNorm codes for medications, and $d_{2}=1,048$ CUI features.
@@
Based on the summary-level co-occurrence matrix, we can derive estimators for CLAIME, CL, and Concate, respectively. To assess the quality of the obtained embeddings, we utilize a benchmark previously introduced in Gan et al., (2023). Specifically, we conduct an evaluation focusing on the embeddings’ capability to identify established relationships among EHR concept pairs. Various categories of known relations, including similarity and relatedness, have been curated from online knowledge sources. Similar pairs of codified concepts were mainly generated based on code hierarchies such as the PheCode hierarchy. Similar pairs of CUIs were extracted from relationships within the online clinical database UMLS. Additionally, we assessed the similarity between mapped code-CUI pairs, leveraging UMLS to map codified concepts to CUIs. For relatedness among CUI-CUI pairs, we extracted information from the UMLS, considering major classes like “may treat or may prevent”, “classifies”, “differential diagnosis”, “method of”, and “causative”. Furthermore, related code-CUI pairs and code-code pairs were curated by mapping disease CUIs to PheCodes, drugs to RxNorm, and procedures to CCS categories. These mapped code pairs were then utilized to evaluate the system’s ability to detect relatedness among codified features or the relatedness between codified and NLP features.
@@
To assess different relationship types, we compute cosine similarities between embedding vectors of related pairs and randomly selected pairs. This approach enables the calculation of the Area Under the Curve (AUC) of the cosine similarities, providing a metric for distinguishing known pairs from random pairs. The random pairs selected have the same semantic type as the known pairs. For example, in investigating the relationship “may treat or may prevent,” we only consider disease-drug pairs. This evaluation strategy allows us to quantify and compare the performance of the EHR embeddings across various relationship types.
@@
Pairs Type Group CLAIME CL Concate Num
CUI-Code Similar CUI-PheCode 0.920 0.920 0.920 345
CUI-RXNORM 0.972 0.919 0.918 49
summary 0.915 0.910 0.909 394
Related May Treat (Prevent) 0.775 0.753 0.753 1302
Classifies 0.905 0.895 0.895 744
ddx 0.773 0.765 0.766 1169
Causative 0.775 0.741 0.742 524
summary 0.800 0.783 0.784 3739
CUI-CUI Similar Parent 0.848 0.855 0.856 422
Sibling 0.796 0.763 0.766 865
summary 0.813 0.794 0.796 1287
Related May Treat (Prevent) 0.767 0.765 0.766 257
Classifies 0.936 0.924 0.925 45
ddx 0.741 0.771 0.773 95
Method_of 0.809 0.796 0.800 15
summary 0.756 0.752 0.752 421
Code-Code Similar PheCode Hierachy 0.914 0.909 0.909 3868
Related ddx 0.790 0.796 0.795 5819
Classifies 0.896 0.900 0.900 4525
May Treat (Prevent) 0.734 0.727 0.725 4598
Causative 0.754 0.756 0.754 2468
summary 0.798 0.799 0.798 17410
@@
Table 1: AUCs of between-vector cosine similarity in detecting known similar or related pairs of codes, CUIs, and codes vs CUIs for embeddings trained by CLAIME, CL, and Concate methods. The last column indicates the number of known pairs.
The results are presented in Table 1. It can be observed that CLAIME excels in embedding relationships across two distinct modalities: code and CUIs. In particular, for the similar CUI-RxNorm group, CLAIME achieves about $6\%$ gain in AUC compared to CL and Concate. It is also much better at capturing causative relationships between codes and CUIs, achieving a $4.4\%$ gain in AUC. Moreover, CLAIME is slightly better in representing the parent-child hierarchical relationships between NLP concepts, with a 2% gain in AUC compared to the other two methods.
Since the data comes from the RA cohort, it is of interest for us to further perform case studies for RA-related codified and CUI concepts. We select $7$ key concepts for detailed analysis: rheumatoid arthritis, C-reactive proteins, stiffness of joint, Reiter’s disease, swollen joint, pneumococcal vaccine, and leflunomide.
@@
For each of the $7$ concepts, we calculate its cosine similarity with all remaining concepts using the embeddings generated by CLAIME, Concate, and CL, respectively. Concepts are ranked by similarity, and a subset comprising the top-$100$ rankings from any one of the three methods is chosen as positive control. Then, we randomly select an equal number of other concepts as negative controls. This process produces an evaluation set for each of the $7$ concepts. Concordance between similarity scores (from CLAIME, Concate, CL) and relevance scores (from GPT3.5 (OpenAI,, 2023), GPT4 (Achiam et al.,, 2023) ) is assessed using Kendall’s tau rank correlation. Relevance scores are obtained by prompting GPT3.5 and GPT4 to rate concept relevance on a scale of 00 to $1$ between the given concept and its respective evaluation set.
Figure 2: Feature-specific Kendall’s ranking correlation between pairs of CLAIME, Concate and CL with GPT3.5 and GPT4 feature scores for 7 concepts, namely, rheumatoid arthritis (RA), C-reactive proteins (CP), stiffness of joint (SoJ), Reiter’s disease (RD), swollen joint (SJ), pneumococcal vaccine (PV), and leflunomide (LF). Note that “GPTx” refers to GPT3.5 on the left and GPT4 on the right, and the lines for “CL-GPTx” overlap with “Concate-GPTx” and hence cannot be seen from the figure.
@@
The results, depicted in Figure 2, reveal that Concate and CL exhibit remarkably similar behavior, with the “CL-GPTx” lines obscured beneath “Concate-GPTx”. CLAIME can effectively encode and preserve relations between these medical concepts and their associated features. Notably, negative rank correlations can be observed for CL-GPT3.5/4 and Concate-GPT3.5/4 regarding “leflunomide” and “pneumococcal vaccine” related codes and CUIs. To gain more insights, some detailed examples are provided in Table 2.
ID Desc Concate CL CLAIME GPT3.5 GPT4
Relatedness or similarity with Leflunomide
PheCode:714.1 Rheumatoid arthritis 0.35 0.35 0.80 0.9 0.9
PheCode:714.2 Juvenile rheumatoid arthritis 0.36 0.37 0.78 0.9 0.8
C0242708 DMARDs 0.36 0.36 0.81 0.9 1.0
RXNORM:5487 Hydrochlorothiazide 0.80 0.81 0.01 0.1 0.2
RXNORM:6809 Metformin 0.79 0.79 0.01 0.2 0.2
RXNORM:2556 Citalopram 0.78 0.78 0.05 0.1 0.1
Relatedness or similarity with Pneumococcal vaccine
C0151735 Injection site reaction NOS 0.29 0.29 0.74 0.8 0.8
CCS:228 Prophylactic vaccinations and inoculations 0.29 0.29 0.64 0.9 1.0
C1445860 Protein antibody 0.04 0.04 0.61 0.6 0.7
RXNORM:10689 Tramadol 0.80 0.80 0.12 0.1 0.1
RXNORM:2556 Citalopram 0.75 0.75 0.15 0.1 0.1
RXNORM:11248 Cyanocobalamin 0.81 0.81 0.18 0.2 0.1
Table 2: Cosine similarity between leflunomide, pneumococcal vaccine and some selected features.
More specifically, leflunomide is a type of disease-modifying antirheumatic drugs (DMARDs) used to treat rheumatoid arthritis by reducing inflammation and permanent damage. Table 2 shows that both Concate and CL fail to capture the close connection between leflunomide, RA, and DMARDs. Additionally, they overestimate similarity with unrelated drugs like hydrochlorothiazide, metformin, and citalopram. In the case of pneumococcal vaccine, it is a vaccine targeting streptococcus pneumoniae infections and can potentially cause injection site reactions. It is related to protein antibodies as they all target the immune system (von Elten et al.,, 2014). Both Concate and CL struggle to capture this similarity and instead overestimate similarity with unrelated drugs like tramadol, cyanocobalamin, and citalopram, used for pain relief, vitamin B12 deficiency, and depressive disorders respectively.

### v2

@@ 4 Simulation Studies
In this section, we evaluate the performance of CLAIME and compare it to CL, Concate, and Kernel PCA (KPCA) through simulation studies. KPCA is a nonlinear PCA method that operates in a high-dimensional feature space induced by a kernel function. Following Gupta et al., (2019), we apply KPCA with Gaussian RBF kernel directly to the co-occurrence matrix $\mathbf{D}$ given that the co-occurrence matrix can be viewed as a generalized $n$-gram similarity matrix. When patient-level data is unavailable, the first three methods can still be applied by performing SVD on various aggregated PMI matrices, as described in Section 2. Since patient-level data is also available in the synthesized datasets, we can additionally apply the gradient descent-based variants: CLAIME-GD and CL-GD, which optimize the objective functions defined in (2.3) and (2.5), respectively. The suffix “-GD” is added to distinguish them from their SVD-based counterparts. To enhance training efficiency, only $10$ negative samples per patient are stochastically sampled. The penalty parameter $\lambda$ is set to $1$, while the learning rate starts at $10^{-4}$ and decays by a factor of $10$ every $10$ epochs to ensure convergence. Both gradient-based algorithms converge when their loss functions change by less than $10^{-6}$.
We generate a patient embedding vector $\mathbf{c}_{i}\in\mathbb{R}^{p}$ for each patient $i$, independently drawn from a multivariate normal distribution $N(\mathbf{0},\mathbf{I}_{p})$. To construct the code and CUI embedding matrices, we first generate $\mathbf{V}^{\star}=\mathbf{R}\mathbf{M}$, where $\mathbf{M}\in\mathbb{R}^{d\times p}$ has i.i.d. standard normal entries, and $\mathbf{R}\in\mathcal{O}_{d,d}$ is an orthonormal matrix obtained via QR decomposition of a random matrix. The rows of $\mathbf{V}^{\star}$ are mean-centered (i.e., each row has zero mean) and rescaled so that its largest singular value is $1$. We then partition $\mathbf{V}^{\star}$ into $\mathbf{V}_{1}^{\star}$ (first $d_{1}$ rows) and $\mathbf{V}_{2}^{\star}$ (remaining $d_{2}$ rows). Performing SVD on $\mathbf{V}_{1}^{\star}$ and $\mathbf{V}_{2}^{\star}$ separately yields the column basis matrices $\mathbf{U}_{1}^{\star}$ and $\mathbf{U}_{2}^{\star}$. Next, we generate the data according to model (2.1), where the patient $i$ is assigned $T_{i}^{(1)}$ codes and $T_{i}^{(2)}$ CUIs, with $T_{i}^{(1)},T_{i}^{(2)}\sim\text{Poisson}(50)$ truncated to exclude zeros. We consider two different cases for the error covariance structure.
@@
For Case 1, we set $\mathbf{\Sigma}^{(M)}=\text{diag}(\mathbf{I}_{p/2},\mathbf{0})/c\in\mathbb{R}^{d_{M}\times d_{M}}$ for $M\in\{1,2\}$, where $c>0$ controls the signal-to-noise ratio (SNR) of the data generating process. For Case 2, we generated $\sigma_{w}\overset{i.i.d.}{\sim}\text{Unif}(0,1)$ and set ${\bm{\Sigma}}_{w,w^{\prime}}^{(M)}=\rho^{|w-w^{\prime}|}\sigma_{w}\sigma_{w^{\prime}}/2$ for $w,w^{\prime}\in[d_{M}]$ and $M\in\{1,2\}$, where $\rho\in(0,1)$ controls both the structure and the SNR. Specifically, larger $\rho$ values yield higher SNR in the synthesized data.
@@
For both cases, we set $d_{1}=d_{2}=d/2$ and conduct three experiments to compare the performance of CLAIME, CLAIME-GD, CL, CL-GD, Concate and KPCA under varying sample sizes $n$, feature sizes $d$, and SNR. Performance is evaluated using the metric $\text{Err}(\widehat{\mathbf{U}},\mathbf{U}^{\star})=\max_{M=1,2}\|\widehat{\mathbf{U}}_{M}\widehat{\mathbf{U}}_{M}^{{\scriptscriptstyle\sf T}}-\mathbf{U}_{M}^{\star}\mathbf{U}_{M}^{\star{\scriptscriptstyle\sf T}}\|_{\mathrm{F}}$, which measures the subspace distance between estimated and true embeddings. We first fix the SNR by setting $c=0.2$ for Case 1 and $\rho=0.8$ for Case 2. In the first experiment, we vary the sample size $n$ in $\{2,4,6\}\times 10^{4}$, while fixing $d=200$ and $p=4$. In the second experiment, we vary $d\in\{100,200,300,400\}$, while fixing $n=2\times 10^{4}$ and $p=4$. For the third experiment, we first fix $n=2\times 10^{4}$, $d=200$ and $p=4$. Then, we vary $c$ equally spaced between $0.2$ and $0.8$ for Case 1, and we vary $\rho$ equally spaced between $0.2$ and $0.8$ for Case 2.
@@
Boxplots of the error metrics $\text{Err}(\widehat{\mathbf{U}},\mathbf{U}^{\star})$, averaged over $100$ repetitions, are presented in Figure 2. The results for CL and CL-GD, CLAIME and CLAIME-GD are very close to each other, as indicated by the proximity of their lines in Figures 2 (a)–(c). This suggests that employing gradient-based methods on patient-level data (i.e., CL-GD and CLAIME-GD) is comparable to utilizing SVD on summary-level PMI matrices (i.e., CL and CLAIME). This observation aligns with the theoretical findings in Propositions 2.1 and 2.2. On one hand, in Figures 2(a) and (b), the errors for CLAIME and CLAIME-GD decrease with an increase in $n$ or a decrease in $d$. This behavior is anticipated and can be theoretically inferred from the upper bounds provided in Theorems 3.13.3. In contrast, the errors for Concate, CL, and CL-GD remain unchanged. This can be theoretically justified by their lower bounds being irrelevant to both $n$ and $d$, as shown in Theorem 3.6. Moreover, it can be observed that KPCA behaves similarly to CL and Concate. From the top panel of Figure 2(c), we observe that all methods incur larger errors as $c$ decreases—that is, as the SNR worsens. Compared to the other methods, CLAIME and CLAIME-GD consistently demonstrate superior performance. This finding supports our initial motivation: the proposed method is more robust in noisier data settings. Similarly, in the bottom panel of Figure 2(c), while other methods exhibit only a slight decline in performance as $\rho$ decreases, CLAIME and CLAIME-GD again significantly outperform the alternatives across all values of $\rho$.
We conducted additional simulation experiments in Supplementary S7.3.1 and S7.3.2 to empirically examine larger $d$ and account for model misspecification. Similar patterns were observed, demonstrating the robustness of our methods across broader settings. Furthermore, in Supplementary S7.3.3, we compared our approach against the Concate and CL methods with debiasing (Remark 3.2), and our method continued to outperform them.
@@ 5 Application to Electronic Health Records Studies
We train a joint representation of codified and narrative CUI features using summary EHR data of patients with at least 1 diagnostic code of Rheumatoid Arthritis (RA) at Mass General Brigham (MGB). The MGB RA EHR cohort consists of 53,716 patients whose longitudinal EHR data have been summarized as co-occurrence counts of EHR concepts within a $30$-day window. We include all EHR diagnosis, medication, and procedure codes which have been rolled up to higher concept levels: diagnostic codes to PheCodes [^1] https://phewascatalog.org/phecodes, procedure codes to clinical classification system (CCS) [^2] https://www.hcup-us.ahrq.gov/toolssoftware/ccs_svcsproc/ccssvcproc.jsp, medication codes to ingredient level RxNorm codes (Liu et al., , 2005). We include all EHR codified features along with CUIs that have occurred more than $15$ times in the cohort. This results in $d=4,668$ features, including $d_{1}=3,477$ codified features with $1,776$ PheCodes for diseases, $238$ CCS codes for procedures, and $1,463$ RxNorm codes for medications, and $d_{2}=1,048$ CUI features.
@@
Based on the summary-level co-occurrence matrix, we can derive estimators for CLAIME, CL, Concate and KPCA, respectively. To assess the quality of the obtained embeddings, we utilize a benchmark previously introduced in Gan et al., (2025). For multimodal learning, we are interested in our method’s ability to capture cross modality relationships. Specifically, we first assess the similarity between mapped code-CUI pairs, leveraging UMLS to map codified concepts to CUIs. For relatedness, we first curate UMLS relationships for CUI-CUI pairs, considering major classes like “may treat or may prevent”, “classifies”, “differential diagnosis”, “method of”, and “causative”. Then, we map disease CUIs to PheCodes, drugs to RxNorm, and procedures to CCS categories to obtain related code-CUI pairs. These mapped code pairs are then utilized to evaluate the system’s ability to detect relatedness between codified and NLP features.
@@
To assess different relationship types, we compute cosine similarities between embedding vectors of known and randomly selected pairs. This enables calculation of the Area Under the Curve (AUC), quantifying how well the embeddings distinguish known pairs from random ones. Random pairs are matched on semantic type; for instance, when evaluating the “may treat or may prevent” relationship, we consider only disease–drug pairs. This strategy allows us to compare embedding performance across relationship types.
@@
Type Group CLAIME CL Concate KPCA Number of known pairs
Similar CUI-PheCode 0.910 0.911 0.911 0.916 345
CUI-RxNorm 0.944 0.898 0.900 0.986 49
Related May Treat (Prevent) 0.775 0.753 0.753 0.711 1302
Classifies 0.905 0.895 0.895 0.887 744
ddx 0.773 0.766 0.765 0.670 1169
Causative 0.775 0.742 0.741 0.695 524
@@
Table 1: AUCs of between-vector cosine similarity in detecting known similar or related pairs of codes vs CUIs for embeddings trained by different methods.
The results in Table 1 show that all methods are more effective at identifying similar pairs than related ones, as indicated by consistently higher AUCs in the “Similar” category. Among the methods, KPCA performs best for detecting similarity, achieving the highest AUCs for both CUI-PheCode and CUI-RxNorm pairs. However, when it comes to the more challenging task of detecting relatedness, the proposed CLAIME method consistently outperforms the baselines, especially KPCA. Notably, CLAIME shows a substantial advantage in identifying causative relationships, with an approximate 3.3% AUC improvement over CL and Concat, and an 8% gain over KPCA. These results underscore CLAIME’s robustness in capturing complex and nuanced associations between medical concepts.
@@
Since the data comes from the RA cohort, it is of interest for us to further perform case studies for RA-related codified and CUI concepts. We select $7$ key concepts for detailed analysis: rheumatoid arthritis (RA), C-reactive proteins (CP), stiffness of joint (SoJ), Reiter’s disease (RD), swollen joint (SJ), pneumococcal vaccine (PV), and leflunomide (LF). For each of the $7$ concepts, we calculate its cosine similarity with all remaining concepts using the embeddings generated by CLAIME, Concate, and CL, respectively. Concepts are ranked by similarity, and a subset comprising the top-$100$ rankings from any one of the three methods is chosen as positive control. Then, we randomly select an equal number of other concepts as negative controls. This process produces an evaluation set for each of the $7$ concepts. Concordance between similarity scores (from CLAIME, Concate, CL) and relevance scores (from GPT3.5 (OpenAI, , 2023), GPT4 (Achiam et al., , 2023) ) is assessed using Kendall’s tau rank correlation. Relevance scores are obtained by prompting GPT3.5 and GPT4 to rate concept relevance on a scale of 0 to $1$ between the given concept and its respective evaluation set. The detailed prompt is given in Supplementary S7.1.
Figure 3: Feature-specific Kendall’s rank correlations between CLAIME, Concate, and CL with GPT-3.5 and GPT-4 scores for 7 concepts. ‘GPTx’ denotes GPT-3.5 (left) and GPT-4 (right); note that CL-GPTx overlaps with Concate-GPTx and is not visible in the figure.
@@
The results, depicted in Figure 3, reveal that Concate and CL exhibit remarkably similar behavior, with the “CL-GPTx” lines obscured beneath “Concate-GPTx”. CLAIME can effectively encode and preserve relations between these medical concepts and their associated features. Notably, negative rank correlations can be observed for CL-GPT3.5/4 and Concate-GPT3.5/4 regarding “leflunomide” and “pneumococcal vaccine” related codes and CUIs. To gain more insights, some detailed examples are provided in Table S1 in Supplementary S7.1.
As shown in Supplementary S6, CLAIME achieved strong performance when partitioning data elements into five modalities: diagnoses, procedures, medications, lab results, and CUIs. Its advantage is most evident in evaluating related relationships, with particularly pronounced improvements for ‘causative,’ ‘classifies,’ and ‘method of’ associations.